OpenStudy (anonymous):

How many solutions are there to the following system of equations? 2x – y = 2 –x + 5y = 3 A. 2 B. 0 C. 1 D. infinitely many

OpenStudy (texaschic101):

we are going to put these in y = mx + b form and compare the slopes and the y intercepts. In y = mx + b form, the slope is in the m position and the y int is in the b position. 2x - y = 2 -y = -2x + 2 y = 2x - 2 the slope is 2 and the y intercept is -2 -x + 5y = 3 5y = x + 3 y = 1/5x + 3/5 the slope is 1/5 and the y intercept is 3/5 Compare : the slopes are different and the y intercepts are different. This means that the line will intersect in 1 place. There is 1 solution. ======================== Learn this : If the slopes are the same, but the y intercepts are different....this means the lines are parallel and have no solution. If the slopes are the same and the y intercepts are the same...this means that the lines coincide and have infinite solutions. if the slopes are different and the y intercepts are different...this means they will intersect giving you 1 solution